Simplify the following fractions leaving the answer in index form.
(2n^2 * 3n^3/6n^7) Can you help me please?

Answer by stanbon(75887)

(2n^2 * 3n^3/6n^7)
= (2*3/6)(n^2*n^3/n^7)
= (1)(1/n^2)
=1/n^2
-----------
Cheers,
Stan H.